OpenStudy (anonymous):

can someone help me please? Find the derivative of the function s(x) = square root x + 3/square root x

OpenStudy (anonymous):

\[\sqrt{x} + 3/\sqrt{x}\]

OpenStudy (anonymous):

you should memorize the derivative of \(\sqrt{x}\) because it is a very common function the derivative of \(\sqrt{x}\) is \(\frac{1}{2\sqrt{x}}\)

OpenStudy (anonymous):

as for \(\frac{3}{\sqrt{x}}\) you can rewrite it as \(3x^{-\frac{1}{2}}\) and use the power rule

OpenStudy (anonymous):

\[1/2\sqrt{x} - 3x^-1/2\]

OpenStudy (anonymous):

is that correct ?^^^

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the first part is right, you still have to take the derivative of \[3x^{-\frac{1}{2}}\] using the power rule

OpenStudy (anonymous):

i just rewrote it in exponential form, you still have to find the derivative

OpenStudy (anonymous):

so -3/2x-1?

OpenStudy (anonymous):

\[\frac{ -3 }{ 2 }x^1\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@satellite73 \[\frac{ 1 }{ 2 }\sqrt{x}+ -3/2x ^{-1}\]

OpenStudy (anonymous):

is that correct?